# Model-agnostic Origin Attribution of Generated Images with Few-shot Examples

### v1

@@ Abstract
Recent progress in visual generative models enables the generation of high-quality images. To prevent the misuse of generated images, it is important to identify the origin model that generates them. In this work, we study the origin attribution of generated images in a practical setting where only a few images generated by a source model are available and the source model cannot be accessed. The goal is to check if a given image is generated by the source model. We first formulate this problem as a few-shot one-class classification task. To solve the task, we propose OCC-CLIP, a CLIP-based framework for few-shot one-class classification, enabling the identification of an image’s source model, even among multiple candidates. Extensive experiments corresponding to various generative models verify the effectiveness of our OCC-CLIP framework. Furthermore, an experiment based on the recently released DALL·E-3 API verifies the real-world applicability of our solution.
@@ 1 Introduction
Recent visual generative models are capable of producing images of exceptional quality, which have raised public concerns regarding Intellectual Property (IP) protection and accountability for misuse. In response to both the challenges and opportunities posed by AIGC, a recent U.S. executive order [3] mandates that all AI-generated content must clearly label its source, such as Stable Diffusion [35]. This makes the attribution of origins for generated images crucial in real-world applications, referring to the process of identifying whether a given image is generated by a particular model.
To address the origin attribution problem above, three main methods have been explored in the community. The first method involves watermarking [46, 26, 32, 48], which requires additional modifications to the generated results, affecting the quality of generation. The second method involves injecting fingerprints [56, 8, 57, 58] into the model during training and employing a supervised classifier to identify these fingerprints. This process necessitates changes in training. Alternation-free approaches have also been proposed, which do not require alterations to the generation or training processes. Specifically, the existing methods utilize inverse engineering [53, 22], based on the idea that a synthetic sample can be most accurately reconstructed by the generator that created it. However, inverse engineering approaches necessitate access to the target model and require sampling many images as references.
@@ 2 Related Work
Origin Attribution of Generated images: Origin attribution, distinct from generated image detection, seeks to determine whether specific images were produced by a particular model. Various strategies have been proposed to address this challenge. One method involves embedding watermarks [46, 26, 32, 48] in images to trace their origins. However, this approach faces limitations as different models might use identical watermarks, which can also be manipulated or removed. Alternatively, injecting unique fingerprints [56, 8, 57, 58] into a model during its training phase enables the identification of these markers using a supervised classifier. Although effective, this technique necessitates modifications to the training process and the model’s architecture. A different, modification-free method is inverse engineering [53, 22], which leverages the principle that a synthetic sample can be most accurately reconstructed by its original generator. However, this approach requires access to the target model and extensive sampling of images for comparison. Furthermore, two studies have explored this area under different settings. One develops a multi-class classifier known as an attributor for fake image attribution. Yet, it is only applicable to text-to-image models and operates within a limited, closed-world scenario[43]. Another study[20] demonstrates a theoretical lower bound on attribution accuracy using smaller datasets (e.g. MNIST [23]). This method is only suitable for unconditional GANs and reveals limited scalability to more complex systems, such as DALL·E-3 [2].
@@ 3.1 Background of CLIP-based Classification
CLIP is a multimodal model pre-trained to predict whether an image matches a text prompt. It includes an image encoder $E_{v}(\cdot)$ and text encoder $E_{t}(\cdot)$. The pre-trained CLIP can perform zero-shot multi-class classification by comparing the image with a list of prompts, each representing a class. Formally, assume we have $K$ classes. Let $X^{v}$$\in\mathcal{X}$ denote an image and $X^{t}_{i}$ represent the $i_{\text{th}}$ prompt representing the $i_{\text{th}}$ class. The predicted $i_{\text{th}}$ class probability of the image $X^{v}$ is computed as follows:
@@
In the classification above, hand-crafted text prompts are applied to represent classes. The prompt designs require specialized knowledge and are time-consuming to create. To alleviate this, Zhou et al. [59] introduced the concept of Context Optimization (CoOp), which employs learnable vectors to refine prompt-related words. Instead of manually designing a prompt, CoOp enables the model to automatically optimize for a suitable prompt. Formally, the prompt for the $i_{\text{th}}$ class can be represented as $X^{p}_{i}=[t]\otimes[CLASS_{i}]$, where $t$ is the learnable context vectors, $\otimes$ is a concatenation operation and $CLASS_{i}$ corresponds to the name of $i_{\text{th}}$ class. The objective of optimization is to minimize the error of predicting ground truth $Y_{i}$ for each $X^{v}_{i}$. This is achieved by using a cross-entropy loss function $\mathcal{L}$ with respect to the learnable prompts. Let $f$ represent the pre-trained CLIP model. The optimization can be described as:
@@ 4.1 Experimental Setting
Evaluation. Each model is evaluated by the Area Under the Receiver Operating Characteristic Curve (AUC). The Receiver Operating Characteristic Curve (ROC) is a graphical representation that plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at various threshold levels. The AUC represents the probability that a classifier will rank a randomly chosen positive instance higher than a randomly chosen negative one. To reduce randomness, each model is trained 10 times with different training sets each time. Then, the mean AUC and the corresponding standard deviation are reported over the test set. A higher AUC score signifies better performance. For each table, the corresponding standard deviations are shown in the supplementary. More experimental details, such as different testing tasks, can be found in the supplementary.
@@ 4.2 Comparison with Baselines
Baselines. Since there are currently no methods perfectly suited to our setting, we conduct a comprehensive evaluation of OCC-CLIP by assessing 12 benchmark methods from various usage areas (see Table 1).
@@
Among the CNN-based and Transformer-based baselines, Swin outshines its counterparts. However, the overall disparity in performance between CNN-based and Transformer-based models is not obvious. The following two methods, Image-Patch and Feature-Patch are both based on ResNet50. Image-Patch performs worse than ResNet50 while Feature-Patch has somewhat but limited improvements compared with ResNet50. This suggests that current methods used to do deepfake detection perform poorly in this few-shot origin attribution scenarios. In addition to the aforementioned classical methods, OCC-CLIP surpasses widely-spread vision-language models such as zero-shot CLIP and CoOp. In detail, OCC-CLIP outperforms zero-shot CLIP by an average of around 31% and CoOp by approximately 2.9%. This progression emphasizes the value of ADA. In a comprehensive assessment, OCC-CLIP emerges as a leading standard, underscoring its exceptional ability in origin attribution.
@@ 4.3 Ablation Study
Sensitivity to Source Models. Table 2 shows the performance of CoOp and OCC-CLIP in origin attribution with eight different source models. During training, the target dataset is from the dataset shown in the leftmost column, and the non-target dataset is COCO. During testing, the target remains the same, but the non-target dataset is replaced with one of the other generated datasets shown in the first row. For example, the source model for the second row in the table is Stable Diffusion [35]. OCC-CLIP demonstrates superior performance in most cases and outperforms the baseline on average in the results from testing with seven other datasets. Overall, our proposed model has exhibited superior performance and generalization capabilities in determining if an image is from the same source model as a set of target images, compared to the baseline.
Table 3: Evaluation of OCC-CLIP on Different Non-target Image Datasets. During training phase, the non-target class is chosen from One of the four datasets (COCO, CC12M, Flickr, and ImageNet). The target dataset is from SD. The performance is averaged across seven different testing tasks.
Methods COCO CC12M Flickr ImageNet
CoOp $0.9263$ $0.8689$ $0.8788$ $0.9377$
OCC-CLIP $\mathbf{0.9548}$ $\mathbf{0.9320}$ $\mathbf{0.9355}$ $\mathbf{0.9710}$
Sensitivity to Selection of Non-target Class. As shown in Table 3, we select non-target images from one of four different clean datasets: COCO, ImageNet, Flickr, or CC12M. The target images are from SD. The average AUC score over 7 different testing tasks is computed. It can be noted that although the choice of the source of non-target images can somewhat affect the performance of origin attribution, our framework consistently outperforms the baseline.
@@
Sensitivity to ADA Settings. We employ four methods for applying ADA: 1) ‘Non-Target’: on half of the non-target images; 2) ‘Target’: on half of the target images; 3) ‘Both’: on half of both the non-target images and target images; 4) ‘T-NT’: on half of the target images which are then treated as non-target ones. The non-target images are selected from the COCO dataset, while the target images are from SD. As shown in Table 4, it is evident that applying ADA on non-target images yields the best performance. This outcome is reasonable, as gradient ascent on the non-target image set enlarges the learned space of this set and further approximates the boundary to the target image set. Applying ADA on ‘T-NT’ is the next most effective method, performing only slightly worse than ‘Non-Target’. However, the boundary learned by ‘T-NT’ may be too tight for effective origin attribution tasks. Applying ADA on ‘Both’ is less effective than doing nothing, as ‘Both’ also enlarges the learned distribution of the target image set. Solely applying ADA on the target set performs significantly worse than doing nothing. This is because this approach not only enlarges the learned distribution space of the target images but also shrinks the learned distribution space of the non-target image set, leading to a higher likelihood of misclassifying many images as target images.
@@ 4.7 Model Attribution with Multiple Source Models
In addition to verifying single models, we also investigate multi-source origin attribution conditions. We explore the process of verifying multiple sources using the eight trained one-class classifiers. Since this involves multi-source origin attribution, the evaluation metric used in this section is accuracy instead of AUC. We evaluate the model under the following conditions: 2 classes (SD, VQ-D), 4 classes (SD, VQ-D, LDM, Glide), 6 classes (SD, VQ-D, LDM, Glide, GALIP, ProGan), and 8 classes (SD, VQ-D, LDM, Glide, GALIP, ProGAN, StyleGAN2, GauGAN). As shown in Table 28, the accuracy decreases with an increasing number of source models, except in the 8 classes condition. We believe this exception for the 8 classes condition occurs because the attribution of each model has been effectively learned by one of the eight classifiers. In every multi-class classification scenario mentioned above, one-class classifiers trained with OCC-CLIP consistently outperform those trained with CoOp.
@@ 5 Conclusions
In this work, we study origin attribution in a practical setting where only a few images generated by a source model are available and the source model cannot be accessed. The introduced problem is first formulated as a few-shot one-classification task. A simple yet effective solution CLIP-based framework is proposed to solve the task. Our experiments on both open-source popular generative models and commercial generation API shows the effectiveness of our framework. Our OCC-CLIP framework can also be applied to solve the few-shot one-classification task in other domains, which we leave in future work. Another future work is to evaluate the adversarial robustness of our framework and build adversarially robust variants.
@@ Appendix 0.C Comparison with Baselines
The standard deviation of each mean of AUC is shown in Table 10 and Table 11. The reason why the CLIP model has a 0 standard deviation is that it is a zero-shot model, which means there are no parameters changed. Therefore, when testing on the same testing set, the result will not change.
@@ Appendix 0.H Sensitivity to Image Processing in Verification Stage
Tables 20 and 21 present a comparative analysis of the performance between OCC-CLIP and baseline methods in response to potential image processing techniques: Gaussian Blur, Gaussian Noise, Grayscale, Rotation, Flip, or a combination of these attacks. The standard deviations and the testing results across seven different tasks, along with their average, are shown. The data indicate that OCC-CLIP exhibits superior robustness compared to the baseline in most cases, demonstrating enhanced resilience to various forms of image manipulation. This underscores the effectiveness of OCC-CLIP in safeguarding against image-based copyright infringements.
@@ Appendix 0.I Sensitivity to Choice of Prompts
We explore the impact of prompt selection on the performance of OCC-CLIP. As indicated in Tables 22 and 23, the choice of prompts influences performance across seven different testing tasks. However, OCC-CLIP consistently outperforms the baseline regardless of the prompt pair chosen.
@@ Appendix 0.J Comparison to Standard Data Augmentation
In a comprehensive evaluation of ADA, we compare its effectiveness with common data augmentation methods: Gaussian Blur, Gaussian Noise, Grayscale, Rotation, Flip, and combinations of these. According to Tables 24 and 25, traditional data augmentation methods generally perform poorly, often degrading performance across many of the seven testing tasks. Only ADA shows improvements, suggesting that standard data augmentation methods may not be suitable for few-shot one-class classification scenarios.
@@ Appendix 0.K Source Model Attribution with Commercial Generation API
To assess OCC-CLIP’s real-world effectiveness, we used the latest commercial digital image-generating model, DALL·E-3 [2], to generate images. As detailed in Section 0.N, 103 prompts were randomly selected from the annotations of COCO’s validation set [24]. In addition to the images generated from these prompts, additional sample images are presented in Figure 5. During the training phase, the target images are those generated by DALL·E-3, while the non-target images are from COCO. In the testing phase, the non-target images are from one of the other generated datasets. The performance of detection across various shot numbers (10, 20, 30, 40, 50) is compared. According to Tables 26 and 27, our method consistently outperforms the baseline across all eight testing tasks and performs well even with as few as 10 shots. Due to limitations in the number of images generated by DALL·E-3, we train the model only once for each testing task.

### v2

# Which Model Generated This Image? A Model-Agnostic Approach for Origin Attribution

@@ Abstract
Recent progress in visual generative models enables the generation of high-quality images. To prevent the misuse of generated images, it is important to identify the origin model that generates them. In this work, we study the origin attribution of generated images in a practical setting where only a few images generated by a source model are available and the source model cannot be accessed. The goal is to check if a given image is generated by the source model. We first formulate this problem as a few-shot one-class classification task. To solve the task, we propose OCC-CLIP, a CLIP-based framework for few-shot one-class classification, enabling the identification of an image’s source model, even among multiple candidates. Extensive experiments corresponding to various generative models verify the effectiveness of our OCC-CLIP framework. Furthermore, an experiment based on the recently released DALL·E-3 API verifies the real-world applicability of our solution. Our source code is available at https://github.com/uwFengyuan/OCC-CLIP.
@@ 1 Introduction
Recent visual generative models are capable of producing images of exceptional quality, which have raised public concerns regarding Intellectual Property (IP) protection and accountability for misuse [15, 31, 34, 33]. In response to both the challenges and opportunities posed by Artificial Intelligence-Generated Content (AIGC), a recent U.S. executive order [3] mandates that all AI-generated content must clearly label its source, such as Stable Diffusion [46]. This makes the attribution of origins for generated images crucial in real-world applications, referring to the process of identifying whether a given image is generated by a particular model.
To address the origin attribution problem above, three main methods have been explored in the community. The first method involves watermarking [57, 37, 43, 59, 35], which requires additional modifications to the generated results, affecting the quality of generation. The second method involves injecting fingerprints [68, 10, 69, 70] into the model during training and employing a supervised classifier to identify these fingerprints. This process necessitates changes in training. Modification-free approaches have also been proposed, which do not require modifications to the generation or training processes. Specifically, the existing methods utilize inverse engineering [64, 29], based on the idea that a synthetic sample can be most accurately reconstructed by the generator that created it. However, inverse engineering approaches necessitate access to the target model and require sampling many images as references.
@@ 2 Related Work
Origin Attribution of Generated images: Origin attribution, distinct from generated image detection, seeks to determine whether specific images were produced by a particular model. Various strategies have been proposed to address this challenge. One method involves embedding watermarks [57, 37, 43, 59] in images to trace their origins. However, this approach faces limitations as different models might use identical watermarks, which can also be manipulated or removed [35]. Alternatively, injecting unique fingerprints [68, 10, 69, 70] into a model during its training phase enables the identification of these markers using a supervised classifier. Although effective, this technique necessitates modifications to the training process and the model’s architecture. A different, modification-free method is inverse engineering [64, 29], which leverages the principle that a synthetic sample can be most accurately reconstructed by its original generator. However, this approach requires access to the target model and extensive sampling of images for comparison. Furthermore, three studies have explored this area under different settings. One develops a multi-class classifier known as an attributor for fake image attribution. Yet, it is only applicable to text-to-image models and operates within a limited, closed-world scenario[54]. One [13] utilizes an iterative multi-step pipeline to detect the origins of images generated by different GANs. However, it requires many samples. The other study[27] demonstrates a theoretical lower bound on attribution accuracy using smaller datasets (e.g. MNIST [30]). This method is only suitable for unconditional GANs and reveals limited scalability to more complex systems, such as DALL·E-3 [2].
@@ 3.1 Background of CLIP-based Classification
CLIP is a multimodal model pre-trained to predict whether an image matches a text prompt. It includes an image encoder $E_{v}(\cdot)$ and text encoder $E_{t}(\cdot)$. The pre-trained CLIP can perform zero-shot multi-class classification by comparing the image with a list of prompts [16], each representing a class. Formally, assume we have $K$ classes. Let $X^{v}$$\in\mathcal{X}$ denote an image and $X^{t}_{i}$ represent the $i_{\text{th}}$ prompt representing the $i_{\text{th}}$ class. The predicted $i_{\text{th}}$ class probability of the image $X^{v}$ is computed as follows:
@@
In the classification above, hand-crafted text prompts are applied to represent classes [16]. The prompt designs require specialized knowledge and are time-consuming to create. To alleviate this, Zhou et al. [71] introduced the concept of Context Optimization (CoOp), which employs learnable vectors to refine prompt-related words. Instead of manually designing a prompt, CoOp enables the model to automatically optimize for a suitable prompt. Formally, the prompt for the $i_{\text{th}}$ class can be represented as $X^{p}_{i}=[t]\otimes[CLASS_{i}]$, where $t$ is the learnable context vectors, $\otimes$ is a concatenation operation and $CLASS_{i}$ corresponds to the name of $i_{\text{th}}$ class. The objective of optimization is to minimize the error of predicting ground truth $Y_{i}$ for each $X^{v}_{i}$. This is achieved by using a cross-entropy loss function $\mathcal{L}$ with respect to the learnable prompts. Let $f$ represent the pre-trained CLIP model. The optimization can be described as:
@@ 4.1 Experimental Setting
Evaluation. Each model is evaluated by the Area Under the Receiver Operating Characteristic Curve (AUC). The Receiver Operating Characteristic Curve (ROC) is a graphical representation that plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at various threshold levels.
The AUC represents the probability that a classifier will rank a randomly chosen positive instance higher than a randomly chosen negative one. To reduce randomness, each model is trained 10 times with different training sets each time. Then, the mean AUC and the corresponding standard deviation are reported over the test set. A higher AUC score signifies better performance. For each table, the corresponding standard deviations are shown in the supplementary. More experimental details, such as different testing tasks and accuracy scores, can be found in the supplementary.
@@ 4.2 Comparison with Baselines
Baselines. Since there are currently no methods perfectly suited to our setting, we conduct a comprehensive evaluation of OCC-CLIP by assessing 12 benchmark methods from various usage areas (see Table 1). The supplementary can find a comparison with other baselines [64, 13].
@@
Among the CNN-based and Transformer-based baselines, Swin outshines its counterparts. However, the overall disparity in performance between CNN-based and Transformer-based models is not obvious. The following two methods, Image-Patch and Feature-Patch are both based on ResNet50. Image-Patch performs worse than ResNet50 while Feature-Patch has somewhat but limited improvements compared with ResNet50. This suggests that current methods used to do deepfake detection perform poorly in this few-shot origin attribution scenarios. In addition to the aforementioned classical methods, OCC-CLIP surpasses widely-spread vision-language models such as zero-shot CLIP and CoOp. In detail, OCC-CLIP outperforms zero-shot CLIP by an average of around 31% and CoOp by approximately 2.9%. This progression emphasizes the value of ADA. In a comprehensive assessment, OCC-CLIP emerges as a leading standard, underscoring its exceptional ability in origin attribution. However, it is harder to distinguish images generated by very similar algorithms trained on similar datasets, such as DeepFloyd, which is based on Stable Diffusion and also trained on the COCO dataset. In this case, our model’s performance does decrease (dropping to $-20.7650\scriptscriptstyle\pm\scriptstyle 4.80e\text{-}$ 2). Nevertheless, our method still outperforms the best baseline (CoOp: $-20.7301\scriptscriptstyle\pm\scriptstyle 7.91e\text{-}$ 2).
@@ 4.3 Ablation Study
Sensitivity to Source Models. Table 2 shows the performance of CoOp and OCC-CLIP in origin attribution with eight different source models. During training, the target dataset is from the dataset shown in the leftmost column, and the non-target dataset is COCO. During testing, the target remains the same, but the non-target dataset is replaced with one of the other generated datasets shown in the first row. For example, the source model for the second row in the table is Stable Diffusion [46].
OCC-CLIP demonstrates superior performance in most cases and outperforms the baseline on average in the results from testing with seven other datasets. Overall, our proposed model has exhibited superior performance and generalization capabilities in determining if an image is from the same source model as a set of target images, compared to the baseline.
Table 3: Evaluation of OCC-CLIP on Different Non-target Image Datasets. During training phase, the non-target class is chosen from one of the four datasets (COCO, CC12M, Flickr, and ImageNet) or from combined datasets. The target dataset is from SD. The performance is averaged across seven different testing tasks.
Methods COCO CC12M Flickr ImageNet Combined
CoOp $0.9263$ $0.8689$ $0.8788$ $0.9377$ $0.9320$
OCC-CLIP $\mathbf{0.9548}$ $\mathbf{0.9320}$ $\mathbf{0.9355}$ $\mathbf{0.9710}$ $\mathbf{0.9651}$
Sensitivity to Selection of Non-target Class. As shown in Table 3, we select non-target images from combined datasets or from one of the four open-domain image datasets: COCO, ImageNet, Flickr, or CC12M. The target images are from SD. The average AUC score over 7 different testing tasks is computed. It can be noted that although the choice of the source of non-target images can somewhat affect the performance of origin attribution, our framework consistently outperforms the baseline.
@@
Sensitivity to ADA Settings. We employ four methods for applying ADA: 1) ‘Non-Target’: on half of the non-target images; 2) ‘Target’: on half of the target images; 3) ‘Both’: on half of both the non-target images and target images; 4) ‘T-NT’: on half of the target images which are then treated as non-target ones. The non-target images are selected from the COCO dataset, while the target images are from SD.
As shown in Table 4, it is evident that applying ADA on non-target images yields the best performance. This outcome is reasonable, as gradient ascent on the non-target image set enlarges the learned space of this set and further approximates the boundary to the target image set. Applying ADA on ‘T-NT’ is the next most effective method, performing only slightly worse than ‘Non-Target’. However, the boundary learned by ‘T-NT’ may be too tight for effective origin attribution tasks. Applying ADA on ‘Both’ is less effective than doing nothing, as ‘Both’ also enlarges the learned distribution of the target image set. Solely applying ADA on the target set performs significantly worse than doing nothing. This is because this approach not only enlarges the learned distribution space of the target images but also shrinks the learned distribution space of the non-target image set, leading to a higher likelihood of misclassifying many images as target images.
@@ 4.7 Model Attribution with Multiple Source Models
In addition to verifying single models, we also investigate multi-source origin attribution conditions. We explore the process of verifying multiple sources using the eight trained one-class classifiers. Since this involves multi-source origin attribution, the evaluation metric used in this section is accuracy instead of AUC. We evaluate the model under the following conditions: 2 classes (SD, VQ-D), 4 classes (SD, VQ-D, LDM, Glide), and 6 classes (SD, VQ-D, LDM, Glide, GALIP, ProGan). As shown in Table 31, the accuracy decreases with an increasing number of source models. In every multi-class classification scenario mentioned above, one-class classifiers trained with OCC-CLIP consistently outperform those trained with CoOp.
@@ 5 Conclusions
In this work, we study origin attribution in a practical setting where only a few images generated by a source model are available and the source model cannot be accessed. The introduced problem is first formulated as a few-shot one-classification task. A simple yet effective solution CLIP-based framework is proposed to solve the task. Our experiments on both open-source popular generative models and commercial generation API shows the effectiveness of our framework. Our OCC-CLIP framework can also be applied to solve the few-shot one-classification task in other domains, which we leave in future work. Another future work is to evaluate the natural and adversarial robustness of our framework [7, 8, 18, 17] and build adversarially robust variants [65, 23].
Acknowledgement: This work is supported by the UKRI grant: Turing AI Fellowship EP/W002981/1, EPSRC/MURI grant: EP/N019474/1. We thank the Royal Academy of Engineering
@@ Appendix 0.C Comparison with Baselines
The standard deviation of each mean of AUC is shown in Table 10 and Table 11. The standard deviation of each mean of Accuracy is shown in Table 12 and Table 13.The reason why the CLIP model has a 0 standard deviation is that it is a zero-shot model, which means there are no parameters changed. Therefore, when testing on the same testing set, the result will not change.
@@ Appendix 0.H Sensitivity to Image Processing in Verification Stage
Tables 22 and 23 present a comparative analysis of the performance between OCC-CLIP and baseline methods in response to potential image processing techniques: Gaussian Blur, Gaussian Noise, Grayscale, Rotation, Flip, or a combination of these attacks. The standard deviations and the testing results across seven different tasks, along with their average, are shown. The data indicate that OCC-CLIP exhibits superior robustness compared to the baseline in most cases, demonstrating enhanced resilience to various forms of image manipulation. This underscores the effectiveness of OCC-CLIP in safeguarding against image-based copyright infringements.
@@ Appendix 0.I Sensitivity to Choice of Prompts
We explore the impact of prompt selection on the performance of OCC-CLIP. As indicated in Tables 24 and 25, the choice of prompts influences performance across seven different testing tasks. However, OCC-CLIP consistently outperforms the baseline regardless of the prompt pair chosen.
@@ Appendix 0.J Comparison to Standard Data Augmentation
In a comprehensive evaluation of ADA, we compare its effectiveness with common data augmentation methods: Gaussian Blur, Gaussian Noise, Grayscale, Rotation, Flip, and combinations of these. According to Tables 26 and 27, traditional data augmentation methods generally perform poorly, often degrading performance across many of the seven testing tasks. Only ADA shows improvements, suggesting that standard data augmentation methods may not be suitable for few-shot one-class classification scenarios.
@@ Appendix 0.K Source Model Attribution with Commercial Generation API
To assess OCC-CLIP’s real-world effectiveness, we used the latest commercial digital image-generating model, DALL·E-3 [2], to generate images. As detailed in Section 0.N, 103 prompts were randomly selected from the annotations of COCO’s validation set [32]. In addition to the images generated from these prompts, additional sample images are presented in Figure 5. During the training phase, the target images are those generated by DALL·E-3, while the non-target images are from COCO. In the testing phase, the non-target images are from one of the other generated datasets. The performance of detection across various shot numbers (10, 20, 30, 40, 50) is compared. According to Tables 28 and 30, our method consistently outperforms the baseline across all eight testing tasks and performs well even with as few as 10 shots. Due to limitations in the number of images generated by DALL·E-3, we train the model only once for each testing task.